Lengthen the two's complement number (both shown in hex) E3 and 3E to 16 bits.

(NOTE: The values represented in each case should remain the same length after lengthening.)

Your 16 bit answer in hex (for E3 in two's complement):

Your 16 bit answer in hex (for 3E in two's complement):

Shorten (if possible) the two's complement numbers (both shown in hex) E3 and 3E to the minimum number of bits needed.
(Note 1: The values represented in each case should remain the same after any shortening)

(Note 2: The minimum number of bits needed for the two numbers need not be the same.)

Your answer in binary (for E3 in two's complement using minimum number of bits):

Your answer in binary (for 3E in two's complement using minimum number of bits):

Will rate, Please show steps clearly, thank you!
Answer #1

Binary representation of E3 using 16 bits = 0000000011100011

1's complement of E3 = 1111111100011100

2's complement of E3 = 1111111100011100 + 1 = 1111111100011101 = FF1D16

Binary representation of 3E using 16 bits = 0000000000111110

1's complement of 3E = 1111111111000001

2's complement of 3E = 1111111111000001 + 1 = 1111111111000010 = FF2C16

2's complement of E3 in minimum number of bits = 100011101

2's complement of 3E in minimum number of bits = 111000010